The reaction between iron(II) (Fe2+) and dichromate (Cr2O2−7) in the presence of a strong acid (H+) is shown. Cr2O2−7+6Fe2++14H+⟶2Cr3++6Fe3++7H2O Determine the volume, in milliliters, of a 0.270 M solution of Mohr's salt ((NH4)2Fe(SO4)2⋅6H2O) needed to completely react with 0.0800 L of 0.270 M potassium dichromate (K2Cr2O7). volume: mL What volume, in milliliters, of 0.270 M K2Cr2O7 is required to completely react with 0.0800 L of a 0.270 M solution of Mohr's salt?
Answer #1

1. Molarity of K2Cr2O7 = 0.270 M = 0.270 mol/L

Volume of K2Cr2O7= 0.0800L

Moles of K2Cr2O7 = molarity * volume = 0.270 mol/L * 0.0800L = 0.0216 mol

From reaction, K2Cr2O7 reacts with 6 moles of mohr's salt

So, the Moles of Mohr's salt = 6*0.0216 mol = 0.1296 mol

Molarity of Mohr's salt = 0.270 M

Volume of Mohr's salt = Moles / Molarity = 0.1296mol/ 0.270M = 0.480L = 480 mL

2. Molarity of Mohr's salt = 0.270 M

Volume of Mohr's salt = 0.0800L

Moles of Mohr's salt = molarity *volume = 0.270 M * 0.0800L = 0.0216 mol

From reaction, K2Cr2O7 reacts with 6 moles of mohr's salt

So, the Moles of K2Cr2O7 = 0.0216 mol/6 = 0.0036 mol

Molarity of K2Cr2O7 = 0.270 M

Volume of of K2Cr2O7 = Moles/ molarity = 0.0036mol / 0.270M = 0.0133L = 13.33 mL